OpenStudy (anonymous):

A car is travelling at 60 km/hr. If the diameter of the tires is 53 cm. Find the rate(angular velocity) in rad/sec and rev/min at which the wheels rotate. With solution please. Thanks. :>

OpenStudy (anonymous):

Pure Rolling Right?

OpenStudy (anonymous):

\[\omega \times r = v \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\omega = \frac{v}{r}\] Try this

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

Plug And Play

OpenStudy (anonymous):

how can i turn it to rad/sec and rev/min

OpenStudy (anonymous):

r is the radius while v is the linear velocity rad/sec * 1rev/2pi rad

OpenStudy (anonymous):

rad 60 sec 1 rev ---- * ----- * -------- sec 1 min 2 pi rad

OpenStudy (anonymous):

30 rev/min

OpenStudy (anonymous):

?

OpenStudy (anonymous):

30/pi rev/min

OpenStudy (anonymous):

why its sec? isn't supposed to be cancelled

OpenStudy (wasiqss):

ishaan u r raisin more n more queries in him mind by doin this way :P simplify it

OpenStudy (anonymous):

ok then bout rad/sec?

OpenStudy (anonymous):

yeah it's supposed to be cancelled i typed that wrong it is \[\frac{30}{\pi} rad/min\] sorry i am making it complicated for you

OpenStudy (anonymous):

no it's okay. :) i appreciate it tho.